# Atomic altermagnetism

### v1

@@ Abstract
In Fig. 3a we show the band structure along a path in the $(k_{x},k_{y})$-plane that bisects the two nodal planes. We identify a spin splitting of 0.2 eV just below the Fermi level. We emphasize that, because Ba2CaOsO6 is a Mott insulator, the bands crossing the Fermi level in Fig. 3a will be pushed away by the on-site repulsion. Further investigation of the impact of strong correlations on the band structure is, however, beyond the scope of our work, which focuses on the symmetry-enforced shape of the spin density on the Os site. In this regard, we note that the $d_{xy}$ form factor of the spin density obtained via DFT+U is consistent with the strong-coupling perspective on the octupolar moment of Os $5d^{2}$, which can be understood as the $z$-component of the pseudo-spin that characterizes the non-Krammers doublet ground-state of the Os atom, arising from the combined effects of crystal fields and SOC on the localized $J=2$ state Paramekanti et al. (2020); Voleti et al. (2021).
@@
Figure 3: DFT+U calculation for Ba2CaOsO6 without (a-c) and with (d-j) SOC . (a), (d) Band structure colored according to the $S_{z}$ spin polarization. (b) Iso-surface of the direct-space spin density (no SOC) around the Os site. (c) Energy iso-surface cut in the plane $k_{z}=0$. Iso-surface of the spin density around the Osmium site (with SOC), projected along the spin components $[100]$ (e), $[010]$ (f) and $[001]$(g). Energy iso-surface cuts in planes that contain the $\Gamma$ point and are normal to directions $[100]$ (h), $[010]$ (i) and $[001]$(j), with color scale representing the spin polarization along the axis normal to the planes.
@@
Once SOC is included, the spin density in Ba2CaOsO6, in agreement with previous reports Fiore Mosca et al. (2022), becomes strongly noncoplanar, with equivalent $d$-wave form factors along the $x$, $y$, and $z$ axes. These results are consistent with the analysis of Ref. Fernandes et al. (2024), which found that the inclusion of SOC makes the spin-density noncoplanar in some altermagnets. Figure 3d shows the band structure (with SOC) along a path in the $x-y$ plane, bisecting the planes normal to $[100]$ and $[010]$ axes. We also show the $z$-component of the spin polarization in a color scale. Note that, in the presence of SOC, the system displays a band gap of around 1 eV. We note, however, that strong correlations will render the material a Mott insulator.
The $x$, $y$ and $z$ components of the spin density around the Os atom are shown in Figs. 3 e-g. Below each spin density, in Figs. 3 h-j, we show energy iso-surface cuts with the spin polarization (component normal to each plane) plotted in color scale. The spin density can be understood as 3 orthogonal $d$-wave form factors corresponding to the $x$, $y$ and $z$ spin components. The nodal structure of the spin density is reflected in the spin polarization of the energy iso-surfaces. In analogy with Eq. 1, a simple model capturing these noncoplanar altermagnetic $d$-wave form factors is given by Fernandes et al. (2024):
$h(\mathbf{k})=\lambda k_{x}k_{y}\sigma_{z}+\lambda k_{z}k_{x}\sigma_{y}+\lambda k_{y}k_{z}\sigma_{x}\text{ .}$ (2)
Note that, because of the cubic symmetry of the crystal structure, and in contrast to La2O3Mn2Se2, all spin components have the same prefactor $\lambda$.
The emergence of the atomic $d$-wave form factors in Ba2CaOsO6, with or without SOC, comes from an interplay between correlations and orbital ordering. The on-site occupation matrix of the $d$ orbitals captures the $d$-wave character Fiore Mosca et al. (2022), which then gets imprinted in both the direct-space spin density and the spin splitting in the band structure.
Figure 4: Estimation of the spin-splitter angle for the non-relativistic calculation of both La2O3Mn2Se2 (a) and Ba2CaOsO6 (b). See in Supplementary Information the calculation of the GMR coefficients.
Spin currents – We calculate the magnitude of the spin currents in the non-relativistic limit for both La2O3Mn2Se2 and Ba2CaOsO6, by quantifying the spin-splitter angle $\alpha$, which measures the efficiency in the charge-to-spin conversion, and the GMR coefficient Šmejkal et al. (2022c). Our results are shown in Fig. 4, where the spin-splitter angle is plotted as a function of the Fermi energy. In the explored energy interval, the spin-splitter angle for Ba2CaOsO6 (Fig. 4a) reaches a maximum of 26 , while for La2O3Mn2Se2 it is even higher, reaching 44. The GMR coefficient (see SI) in La2O3Mn2Se2 reaches 900% for energies close to the Fermi energy, which is one order of magnitude higher than the 100% scale of the initial altermagnetic GMR prediction, considering RuO2 Šmejkal et al. (2022c). While La2O3Mn2Se2 is not metallic, we expect the spin-splitter angle and the GMR coefficient to be equally large in related metallic altermagnetic candidates on layered Lieb lattices, such as UCr2Si2Lemoine et al. (2018), KV2Se2O Jiang et al. or RbV2Se2O Zhang et al. (2024). In addition to the $d$-wave form factor in the sites hosting the magnetic dipoles, the $d$-wave atomic spin density will be present, without dipole, at the O site in RbV2Se2O and KV2Se2O, and at the U site at UCr2Si2C. In the SI, we show calculations of the electronic structure for UCr2Si2C, and we comment as well on the nature of the spin conductivity tensor in Ba2CaOsO6(with SOC), when the spin density is described by three noncoplanar $d$-wave form factors.
Discussion – While the notion of multipolar magnetism is naturally connected with altermagnetism Bhowal and Spaldin (2024); Fernandes et al. (2024), one should be careful when comparing these two concepts. Not every direct-space ferroically ordered magnetic multipole, corresponds to the collinear $d$-, $g$-, or $i$-wave altermagnetic order with the characteristic anisotropic nodal spin polarization in the momentum-space electronic structure. As an illustrative example, consider Mn3Sn, which has a phase between $T_{N2}=50$K and $T_{N1}=420$ K in which Mn magnetic moments form an inverse triangular spin structure. Although the material exhibits octupolar order $T_{x}^{\gamma}$ Suzuki et al. (2017), its magnetic order on the crystal is non-collinear and, therefore, is not altermagnetic. Correspondingly, the resulting spin polarization in the momentum space takes noncollinear and nodeless form Hellenes et al. (2023).
Conclusions— In this work we extend the known mechanisms that induce altermagnetism, to include pure atomic altermagnetism. The interplay between strong correlations and orbital occupation (possibly with strong SOC) generates a spin density which lacks magnetic dipole moments, while having a $d$-wave component which is also reflected in the corresponding spin-polarization symmetry of the electronic structure in the momentum space. In La2O3Mn2Se2 we demonstrate the coexistence of dipole and higher-order partial-wave components on the Mn-sites, with the absence of the dipole and the presence of the higher-order partial-wave component (atomic altermagnetism) on the O-sites. We estimate the spin-splitter angle and the GMR coefficient in this material as 44 degrees and 900%, respectively.
We identified pure atomic altermagnetism with $d$-wave symmetry in the non-relativistic limit of Ba2CaOsO6. When including SOC, the spin density has the structure of three noncoplanar (and orthogonal) $d$-waves form factors. In all cases, we identify one-to-one correspondence between the ferroically ordered $d$-wave harmonic of the spin density in direct space, and the angular dependence of the spin splitting around the $\Gamma$ point.
@@
Acknowledgments— RJU, VKB, RZ, JS and LS acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)- TRR 173– 268565370 (project A03) and TRR 288– 422213477 (project A09 and B05). TJ acknowledges Ministry of Education of the Czech Republic Grant CZ.02.01.01/00/22008/0004594 and ERC Advanced Grant no. 101095925. R.M.F. was supported by the Air Force Office of Scientific Research under Award No. FA9559-21-1-0423.

### v2

@@ Abstract
In Fig. 3a we show the band structure along a path in the $k_{z}=0$ plane that bisects the two nodal planes. We identify a spin splitting of 0.2 eV just below the Fermi level. We emphasize that, because Ba2CaOsO6 is a Mott insulator, bands crossing the Fermi level in Fig. 3a will be pushed away by the on-site repulsion. A full treatment of strong correlations is beyond the scope of this work, which focuses on the symmetry-enforced shape of the spin density. In this regard, we note that the $d_{xy}$ form factor of the spin density obtained via DFT+U is consistent with the strong-coupling perspective on the octupolar moment of Os $5d^{2}$, which can be understood as the $z$-component of the pseudo-spin that characterizes the non-Krammers doublet ground-state of the Os atom, arising from the combined effects of crystal fields and SOC on the localized $J=2$ state Paramekanti et al. (2020); Voleti et al. (2021).
@@
Figure 3: DFT+U results for Ba2CaOsO6 without (a–c) and with (d–j) SOC. (a,d) Band structures colored by $S_{z}$ spin polarization. (b) Iso-surface of the direct-space spin density (no SOC) around the Os site. (c) Energy iso-surface in the $k_{z}=0$ plane. (e–g) Spin density iso-surfaces with SOC, projected along $[100]$, $[010]$, and $[001]$. (h–j) Energy iso-surface cuts in planes normal to $[100]$, $[010]$, and $[001]$, colored by spin polarization normal to each plane.
@@
Upon inclusion of SOC, the spin density in Ba2CaOsO6 becomes strongly noncoplanar, with equivalent $d$-wave form factors along $x$, $y$, and $z$, consistent with Refs. Fiore Mosca et al., 2022 and Fernandes et al., 2024. Figure 3d shows the band structure (with SOC) along a path in the xyx-y plane, bisecting the planes normal to $[100]$ and $[010]$ axes. The $z$-component of the spin polarization is shown in color scale. SOC opens a  1 eV band gap. We note, however, that strong correlations will make the system a Mott insulator.
In analogy with Eq. 1, a simple model capturing these noncoplanar altermagnetic $d$-wave form factors is given by Fernandes et al. (2024):
@@
Note that, because of the cubic symmetry of the crystal structure, and in contrast to KV2Se2O, all spin components have the same prefactor $\lambda$.
The DFT spin density projected along $x$, $y$, and $z$ is shown in Figs. 3e–g. Below, Figs. 3h–j show energy iso-surfaces with spin polarization (component normal to the plane) in color. Comparing the DFT calculations with Eq. 2, we can understand the spin density as 3 orthogonal $d$-wave form factors corresponding to the $x$, $y$ and $z$ spin components. The nodal structure of the spin density is reflected in the spin polarization of the energy iso-surfaces.
The $d$-wave form factors in Ba2CaOsO6 arise from an interplay between correlations and orbital order. The on-site occupation matrix of the $d$ orbitals captures the $d$-wave character Fiore Mosca et al. (2022), which then gets imprinted in both the direct-space spin density and the spin splitting in the band structure. A similar orbital-order mechanism was recently studied in a Hubbard square lattice model Giuli et al. (2025).
Figure 4: Spin-splitter angle from non-relativistic DFT for KV2Se2O (a) and Ba2CaOsO6 (b).
Spin currents – We compute spin currents in the non-relativistic limit for KV2Se2O and Ba2CaOsO6. We quantify the spin-splitter angle $\alpha$ González-Hernández et al. (2021) and the GMR coefficient Šmejkal et al. (2022c). Figure 4 shows $\alpha$ as a function of Fermi energy, peaking at 26 for Ba2CaOsO6 and 42 for KV2Se2O. The GMR coefficient in KV2Se2O reaches 800% (see SI), nearly an order of magnitude above original prediction in RuO2 Šmejkal et al. (2022c). Comparable values are expected in related metallic Lieb-lattice altermagnets such as RbV2Se2O Zhang et al. (2025). In the SI, we comment on the nature of the spin conductivity tensor in Ba2CaOsO6  when the spin density is described by three noncoplanar $d$-wave form factors.
Discussion—While the notion of multipolar magnetism is naturally connected with altermagnetism Bhowal and Spaldin (2024); Fernandes et al. (2024), one should be careful when comparing these two concepts. Not every direct-space ferroically ordered magnetic multipole, corresponds to the collinear $d$-, $g$-, or $i$-wave altermagnetic order with the characteristic anisotropic nodal spin polarization in the momentum-space electronic structure. As an illustrative example, consider Mn3Sn, which has a phase between $T_{N2}=50$K and $T_{N1}=420$ K in which Mn magnetic moments form an inverse triangular spin structure. Although the material exhibits octupolar order $T_{x}^{\gamma}$ Suzuki et al. (2017), its magnetic order on the crystal is non-collinear and, therefore, is not altermagnetic. Correspondingly, the resulting spin polarization in the momentum space takes noncollinear and nodeless form Hellenes et al. (2023).
Conclusions— In this work we extend the known mechanisms that induce altermagnetism, to include pure atomic altermagnetism. The interplay between strong correlations and orbital occupation (possibly with strong SOC) generates a spin density which lacks magnetic dipole moments, while having a $d$-wave component which is also reflected in the corresponding spin-polarization symmetry of the electronic structure in the momentum space. We identified pure atomic altermagnetism with $d$-wave symmetry in the non-relativistic limit of Ba2CaOsO6. When including SOC, the spin density has the structure of three noncoplanar (and orthogonal) $d$-wave form factors. In all cases, we identify one-to-one correspondence between the ferroically ordered $d$-wave harmonic of the spin density in direct space, and the angular dependence of the spin splitting around the $\Gamma$ point.
In a family of compounds in the Lieb lattice (e.g. KV2Se2O) we demonstrate the coexistence of dipole and higher-order partial-wave components on the V-sites, with the absence of the dipole and the presence of the higher-order partial-wave component (atomic altermagnetism) on the O-sites. We estimate the spin-splitter angle and the GMR coefficient in this material as 42 degrees and 800%, respectively.
This work thus extends the search for altermagnetic candidates to materials that, while breaking time-reversal symmetry, do not have dipolar order. The detection of the magnetic order without local magnetic dipoles might represent an intriguing challenge for the conventional techniques like neutron scattering. The spin-polarized electronic structure, and the spintronic responses associated with the altermagnetic order, can serve as experimental probes of these higher-order partial-wave forms of magnetic ordering. Besides Ba2CaOsO6, there is a range of materials that have been previously identified as candidates for magnetic octupolar order that may deserve further investigation, in the context of the higher-partial-wave ordering and atomic altermagnetism Santini and Amoretti (2000); Lovesey et al. (2003); Aoki et al. (2008); Matsumura et al. (2009); Sibille et al. (2020), including the metallic cage compound PrV2Al20, which has recently drawn significant attention Tsujimoto et al. (2014); Freyer et al. (2018); Patri et al. (2019); Sorensen and Fisher (2021); Ye et al. (2024).
@@
Acknowledgments—RJU, VKB, RZ, JS and LS acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)- TRR 173– 268565370 (project A03) and TRR 288– 422213477 (project A09 and B05). TJ acknowledges Ministry of Education of the Czech Republic Grant CZ.02.01.01/00/22008/0004594 and ERC Advanced Grant no. 101095925. R.M.F. was supported by the Air Force Office of Scientific Research under Award No. FA9559-21-1-0423. The authors acknowledge the computing time granted on the supercomputer Mogon at Johannes Gutenberg University Mainz (hpc.uni-mainz.de).